OpenStudy (anonymous):

a car accelerates uniformly from rest to a speed of 23.7 m/s in 6.5. find the distance the car travels during this time?

OpenStudy (anonymous):

where a= acceleration, v= final velocity, u=initial velocity, and d=distance: v=23.7m/s u=0 first we need to find acceleration: a=(v-u)/t a=(23.7)/6.5s a=3.6461538 m/s2 then using the formula d=ut+1/2at^2, we can find the distance: d=1/2(3.6461538)(42.25) d=154.04999805